OpenStudy (anonymous):

I have a question on alg 2

OpenStudy (anonymous):

IM in algebra two...ill try to help but i apologize if i cant

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt[3]{6^{-6}} }\]

OpenStudy (anonymous):

It wants me to simplify in radical form and justify each step by identifying the properties of rational exponents used

OpenStudy (anonymous):

do u have any choices?

OpenStudy (phi):

I would change \[ \sqrt[3]{stuff} = stuff^\frac{1}{3} \]

OpenStudy (anonymous):

just a minute. ive never seen this before but i think i might be able to try to solve it?

OpenStudy (anonymous):

unless there are some answers idk. im sorry. the answer i got is probably wayyyyyy offf i think i got  if thats not right im sorry. i have to go now. bye

OpenStudy (anonymous):

Alright thanks haha

OpenStudy (phi):

one way to do this is first convert from "radical" form to "exponent" form using the rule \[ \sqrt[a]{b}= b^\frac{1}{a} \] You would re-write \[ \frac{ 1 }{ \sqrt[3]{6^{-6}} } = \frac{ 1 }{\left(6^{-6}\right)^\frac{1}{3}} \] if we concentrate just on the bottom, we can use this rule \[ \left(a^b \right)^c = a^{b\cdot c} \] to write \[ \left(6^{-6}\right)^\frac{1}{3} = 6^{-6\cdot \frac{1}{3}}= 6^{-2}\] and you now have \[ \frac{ 1 }{6^{-2} }\] there is another rule: you can "flip" the fraction using this: \[ a^{-b} = \frac{1}{ a^{b}} \] or \[ a^b= \frac{1}{ a^{-b}} \] in other words, \[ \frac{ 1 }{6^{-2} } = 6^2\] Finally \[ 6^2 = 36\]